[Hallc_running] [New Logentry] RC update - July 25,2025
arunin at jlab.org
arunin at jlab.org
Fri Jul 25 11:15:02 EDT 2025
Logentry Text:
--
Data taking proceeded smoothly over the past 24 hours. We began with the z = 0.5 π-SIDIS setting and, by the end of the owl shift, we were wrapping up this kinematic configuration.
This morning, we encountered an issue with the target cryogenics system and we contacted Chandan(https://logbooks.jlab.org/entry/4426588, https://logbooks.jlab.org/entry/4426591, https://logbooks.jlab.org/entry/4426600, https://logbooks.jlab.org/entry/4426617, https://logbooks.jlab.org/entry/4426662). We paused data taking until the targets cooled down again. During this downtime, we requested a beam current of 50 μA and changed the polarity of the SHMS magnets (https://logbooks.jlab.org/entry/4426601).
We have now started running with the z = 0.90 π+ SIDIS setting for LH2.
We will hold our regular RC meeting today at 2:30 PM. The Zoom link is provided below.
https://jlab-org.zoomgov.com/j/1618227115?pwd=8JkhFg6YNesVc3HNgWjwUSIKIu...
Meeting ID: 161 822 7115
Passcode: 253463
---
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/4426673
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20250725/8b7c23eb/attachment.htm>
More information about the Hallc_running
mailing list